OpenStudy (anonymous):

an open rectangular box is to be constructed from a piece of cardboard, by cutting squares of length x from each corner and folding up the sides. a= 25 cm b= 46 cm what is the maximum volume of such a box ?

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

oh maybe it wasn't you the first time

OpenStudy (anonymous):

no , it was my friend

OpenStudy (anonymous):

any ideas on how to do it ?

OpenStudy (anonymous):

volume is \[v(x)=x(25-2x)(46-2x)=4 x^3-142 x^2+1150 x\] derivative is \[v'(x)=12 x^2-284 x+1150\] set = 0 and solve to get \[x = \frac{71-\sqrt{1591}}{6}\]

OpenStudy (anonymous):

probably want the decimal it is about 5.19

OpenStudy (anonymous):

if it is not clear let me know

OpenStudy (anonymous):

thanks a lot , very clear now

OpenStudy (anonymous):

yw